OpenStudy (anonymous):

Let vector F = (15a^2x + 5ay^2) i + (5z^3 − 12ay) j − (3z + 5x^2 + 5y^2) k . (a) Find the values of a such that div F = 0 a = (smaller value) a = (larger value) (b) Find the value of a such that div F is a minimum a =

OpenStudy (anonymous):

can you help me out @dan815 please

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\[\large \mathrm{div}\left( P\hat{i},~ Q\hat{j},~ R\hat{k}\right) = P_x + Q_y + R_z\]

ganeshie8 (ganeshie8):

find the divergece set it equal to 0 and solve \(a\)

OpenStudy (anonymous):

how would I find the minimum?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

is the divergence 3(5a^2-4a-1)

ganeshie8 (ganeshie8):

whats the minimum value of a parabola ?

OpenStudy (anonymous):

I got a=-1/5 and a=1

ganeshie8 (ganeshie8):

you don't need x intercepts

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

so how do i find the value of "a" such that divF is a minimum?

ganeshie8 (ganeshie8):

wolfram says minimum = -27/5 http://www.wolframalpha.com/input/?i=minimum+3%285a%5E2-4a-1%29+

ganeshie8 (ganeshie8):

Oh you want the value of \(a\) that minimizes the divergence

ganeshie8 (ganeshie8):

then just take average value of xintercepts

OpenStudy (anonymous):

so 2/5?

ganeshie8 (ganeshie8):

Yep!

OpenStudy (anonymous):

so is there a easy way to find that value?

OpenStudy (anonymous):

or the best thing to use is wolframalpha? haha

OpenStudy (anonymous):

thanks :)